# FinEval-KR: A Financial Domain Evaluation Framework for Large Language Models’ Knowledge and Reasoning

### v1

@@ 1 Introduction
In recent years, rapid LLM development has led the transformation of artificial intelligence. LLMs demonstrate strong natural language processing capabilities and inspire application innovations across various fields, including scientific research Zhang et al. (2024), financial services Nie et al. (2024b), content creation Betker et al. (2023), etc. However, achieving satisfactory performance for complex tasks, such as financial decision making, proves difficult when relying solely on knowledge introduced during training or instructions Wang and Brorsson (2025); Liu et al. (2025); Wang et al. (2024). Reasoning ability, that is, the ability of logical deduction, problem solving, and abstract thought, stands as a core hallmark of advanced intelligence and becomes crucial for assessing LLM intelligibility and applicability Wang and Song (2024); Li et al. (2024c); Valmeekam et al. (2023b). Therefore, it is critical to accurately assess the knowledge and reasoning abilities of LLMs. This helps to understand the shortcomings of the model to support targeted optimization.
@@
Insufficient capability decoupling. Mainstream benchmarks typically evaluate a model’s capabilities based on its performance across various tasks. Their performance represents either the model’s knowledge capacity Nie et al. (2024a); Liu and Jin (2024) or its reasoning ability Saparov and He (2023); Geva et al. (2021). However, our experiments demonstrate that LLM performance in reasoning tasks is influenced by both knowledge and reasoning ability. Therefore, it is essential to decouple and quantify them separately to achieve a more accurate capability characterization.
@@
Neglect of the cognitive science perspective in financial LLM benchmark. Turning the perspective to the financial domain, a majority of work concentrates on assessing model performance in financial NLP tasks and knowledge coverage, with only a few focusing the reasoning ability. Even so, these methods lack systematic design from a cognitive science perspective, making it challenging to uncover differences in model capabilities across various cognitive levels.
@@
To address these questions, this paper first conducts a preliminary experiment taking the financial domain as an entry point to empirically verify the fundamental role of knowledge in domain reasoning tasks (see Appendix A for details). Furthermore, to bridge the gap in relevant research and provide reproducible research resources, the paper makes the following main contributions:
  • A novel evaluation framework. We design a novel LLM evaluation framework that, for the first time, effectively separates domain knowledge assessment and reasoning assessment from a sole task performance metric, and proposes novel metrics: knowledge score and reasoning score. Moreover, inspired by cognitive science, a prerequisite for an LLM to solve complex reasoning problems is that it possesses a certain level of “cognitive abilities”. This motivates us to propose another new metric, cognitive score, based on Bloom’s cognitive taxonomy, to explore in a fine-grained manner the abilities exhibited by LLMs when facing reasoning tasks.
  • A new open-sourced financial domain reasoning dataset [^1] The dataset will be open-sourced after acceptance of the paper.. The released dataset encompasses 22 key financial subfields and provides sample-level knowledge labels, step-level reasoning processes, and cognitive abilities required for the tasks. The dataset contributes to the community by providing a comprehensive and specialized benchmark for research in financial and cognitive reasoning.
  • Evaluation results of mainstream LLMs. Based on the proposed framework and dataset, we conduct a comprehensive evaluation of the current mainstream LLMs, which verifies the effectiveness of the proposed evaluation methodology and yields a series of insightful conclusions (see Section 4.3 and Appendix G for details).

### v3

@@ 1 Introduction
In recent years, rapid LLM development has led the transformation of artificial intelligence. LLMs demonstrate strong natural language processing capabilities and inspire application innovations across various fields, including scientific research zhang2024comprehensive, financial services nie2024survey, content creation betker2023improving, etc. However, achieving satisfactory performance for complex tasks, such as financial decision making, proves difficult when relying solely on knowledge introduced during training or instructions wang2025can; liu2025fin; wang2024re. Reasoning ability, that is, the ability of logical deduction, problem solving, and abstract thought, stands as a core hallmark of advanced intelligence and becomes crucial for assessing LLM intelligibility and applicability wang2024mars; li2024llms; valmeekam2023planbench. Therefore, it is critical to accurately assess the knowledge and reasoning abilities of LLMs. This helps to understand the shortcomings of the model to support targeted optimization.
@@
Insufficient capability decoupling. Mainstream benchmarks typically evaluate a model’s capabilities based on its performance across various tasks. Their performance represents either the model’s knowledge capacity nie2024cfinbench; liu2024mtfineval or its reasoning ability DBLP:conf/iclr/Saparov023; geva2021did. However, our experiments demonstrate that LLM performance in reasoning tasks is influenced by both knowledge and reasoning ability. Therefore, it is essential to decouple and quantify them separately to achieve a more accurate capability characterization.
@@
Neglect of the cognitive science perspective in financial LLM benchmark. While some benchmarks have begun to assess reasoning ability, they generally lack a design grounded in cognitive science–a critical omission for the financial domain. Unlike the deductive reasoning of law or the diagnostic processes of medicine, the essence of financial decision-making is a quantitative game against future uncertainty. This requires a spectrum of higher-order cognitive abilities that extend far beyond simple knowledge application zhang2025xfinbench. For instance, evaluating the impact of a central bank’s interest rate policy requires not only applying knowledge of rate-to-exchange dynamics but also evaluating the complex interplay between market sentiment and strategic policy expectations. This distinction underscores the need for a framework like Bloom’s Taxonomy as an essential diagnostic tool to pinpoint the specific cognitive deficiencies that hinder advanced financial reasoning in LLMs.
@@
We begin by empirically verifying the fundamental role of knowledge in domain-specific reasoning through a preliminary experiment in the finance sector. Building on these initial findings, we make the following contributions:
  • A novel evaluation framework. We propose a novel LLM evaluation framework that disentangles the assessment of domain knowledge and reasoning ability from task performance metrics. This allows us to introduce two distinct metrics: a Knowledge Score and a Reasoning Score. Furthermore, inspired by cognitive science, we posit that complex reasoning relies on a hierarchy of cognitive abilities. This motivates our third metric, the Cognitive Score, which leverages Bloom’s Taxonomy to provide a fine-grained analysis of the cognitive processes employed by LLMs during reasoning.
  • A new open-source Chinese-language dataset for financial reasoning [^1] https://github.com/SUFE-AIFLM-Lab/FinEval-KR. The released dataset encompasses 22 key financial subfields, its primary contribution lies in its multi-layered annotations, where each sample includes knowledge point labels, step-by-step reasoning chains, and the required cognitive skills. As a comprehensive and deeply annotated resource, it is designed to serve as a specialized benchmark to advance research in both financial and cognitive reasoning.
  • Evaluation results of mainstream LLMs. Based on the proposed framework and dataset, we conduct a comprehensive evaluation of the current mainstream LLMs, which verifies the effectiveness of the proposed evaluation methodology and yields a series of insightful conclusions (see Section 5.3 and Appendix I for details).